# Spectral finiteness, quantum norm continuity and classical points

### 13 Mar 2026 v2

@@ 1 Quantum uniform continuity in several guises, and consequences
  1. Claim: the unit of any of the semigroups $\mathcal{S}$}, $\mathcal{K}$} or ${\mathcal{C}}({\mathbb{G}})^{*}$, if it exists, must be the counit. That the variants of the claim are mutually equivalent follows from the fact that arbitrary continuous functionals are expressible as linear combinations of at most four states (e.g. by [28, Theorem III.4.2(ii)]). Having thus reduced the problem to its ${\mathcal{C}}({\mathbb{G}})^{*}$ variant, write $\psi$ for the assumed unit of that convolution monoid and
@@
    The $\varepsilon_{{\mathcal{F}}}$ being restrictions to ${\mathcal{O}}({\mathbb{G}})$ of continuous functionals (i.e. members of ${\mathcal{C}}({\mathbb{G}})^{*}$) by [34, (5.37) and surrounding discussion],
@@
    for $\mathcal{K}$}: this follows from the identification [5, Proposition 7.1.5]
@@
  • The word length function [31, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ (typically assumed symmetric, i.e. with the dual $\alpha^{*}\in{\mathcal{F}}$ whenever $\alpha\in{\mathcal{F}}$) is
@@
  • $\Gamma$ has the property of rapid decay (RD) [31, Proposition and Definition 3.5] (with respect to $\ell_{{\mathcal{F}}}$) if
@@
The following result shows (e.g. via Example˜1.10) that there do exist uniform≤1 representations with infinite spectrum.
@@
Observe first that the individual positive operators $F_{\alpha_{n}}$ of ˜1-1 are identities: [31, Proposition 4.7] proves this globally, assuming the quantum group as a whole has property RD, but the argument (relying on [31, Lemmas 4.3 and 4.6]) goes through in the local version needed here.
Write $H:=\bigoplus_{n}V_{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question. We have to argue, per Corollary˜1.7, that for $v\in H$ with
@@
($\operatorname{\mathrm{supp}}$ defined as in ˜1-8) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large. The unimodularity condition $F_{\alpha}=1$ ensures [34, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
@@
  • the free orthogonal group ${\mathbb{G}}:={\mathbb{O}}^{+}(n)$, $n\geq 3$ (underlying function algebra ${\mathcal{C}}({\mathbb{G}})=A_{o}(I_{n})$ in [31, §4.3]), with ${\mathcal{O}}({\mathbb{G}})$ freely generated by a unitary $n\times n$ matrix $u^{\alpha}=\left(u^{\alpha}_{ij}\right)$, $1\leq i,j\leq n$ of self-adjoint elements;
@@
will do by [31, Theorems 4.9 and 4.10] respectively. $\blacklozenge$

### 20 Mar 2026 v3

@@ 1 Quantum uniform continuity in several guises, and consequences
  1. Claim: the unit of any of the semigroups $\mathcal{S}$}, $\mathcal{K}$} or ${\mathcal{C}}({\mathbb{G}})^{*}$, if it exists, must be the counit. That the variants of the claim are mutually equivalent follows from the fact that arbitrary continuous functionals are expressible as linear combinations of at most four states (e.g. by [29, Theorem III.4.2(ii)]). Having thus reduced the problem to its ${\mathcal{C}}({\mathbb{G}})^{*}$ variant, write $\psi$ for the assumed unit of that convolution monoid and
@@
    The $\varepsilon_{{\mathcal{F}}}$ being restrictions to ${\mathcal{O}}({\mathbb{G}})$ of continuous functionals (i.e. members of ${\mathcal{C}}({\mathbb{G}})^{*}$) by [35, (5.37) and surrounding discussion],
@@
    for $\mathcal{K}$}: this follows from the identification [6, Proposition 7.1.5]
@@
  • The word length function [32, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ (typically assumed symmetric, i.e. with the dual $\alpha^{*}\in{\mathcal{F}}$ whenever $\alpha\in{\mathcal{F}}$) is
@@
  • $\Gamma$ has the property of rapid decay (RD) [32, Proposition and Definition 3.5] (with respect to $\ell_{{\mathcal{F}}}$) if
@@
Theorem˜1.9 shows (e.g. via Example˜1.10) that there do exist uniform≤1 representations with infinite spectrum; an auxiliary remark precedes it.
For $v\in H$ we write
Lemma 1.8.
A unitary representation $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ of a compact quantum group $\mathbb{G}$} is uniform≤1 as a Banach-space representation provided
Proof.
Per Lemma˜1.4, we have to verify Proposition˜1.6’s condition ˜a. A slightly more careful reworking of the implication ˜dr $\Rightarrow$ ˜a (building on the proof of [31, Proposition 2.2]) will deliver that outcome given the present hypotheses.
Recall to that end that said proof begins by assuming a weak-0-convergent net $\left(\varphi_{\lambda}\right)_{\lambda}\subset{\mathcal{C}}({\mathbb{G}})^{*}$ with $U_{\varphi_{\lambda}}:=\left(\varphi_{\lambda}\otimes\operatorname{id}\right)U$ norm-bounded away from 0 and selecting corresponding
$\psi_{\lambda}:=\omega_{v_{\lambda},w_{\lambda}}\in{\mathcal{L}}(H)_{*},\quad v_{\lambda},w_{\lambda}\in H_{\leq 1}$
with $\left|\psi_{\lambda}U_{\varphi_{\lambda}}\right|$ sufficiently close to $\left\|U_{\varphi_{\lambda}}\right\|$. Write
$P^{{\mathcal{F}}}:=\sum_{\alpha\in{\mathcal{F}}}P^{\alpha},\quad{\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$
The hypothesis ensures that for finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ sufficiently large,
$\left|\psi^{{\mathcal{F}}}_{\lambda}U_{\varphi_{\lambda}}-\psi_{\lambda}U_{\varphi_{\lambda}}\right|$ $=\left|\varphi_{\lambda}\left(\operatorname{id}\otimes\psi^{{\mathcal{F}}}_{\lambda}\right)U-\varphi_{\lambda}\left(\operatorname{id}\otimes\psi_{\lambda}\right)U\right|$
$\psi^{{\mathcal{F}}}_{\lambda}$ $\in\left\{\omega_{P^{{\mathcal{F}}}v_{\lambda},w_{\lambda}},\ \omega_{v_{\lambda},P^{{\mathcal{F}}}w_{\lambda}},\ \omega_{P^{{\mathcal{F}}}v_{\lambda},P^{{\mathcal{F}}}w_{\lambda}}\right\}$
is $\lambda$-uniformly small, so we may assume all $v_{\lambda}$ and $w_{\lambda}$ finitely supported on a single ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$. The truncation $\left(1\otimes P^{{\mathcal{F}}}\right)U$ belonging to the algebraic tensor product
${\mathcal{C}}({\mathbb{G}})^{{\mathcal{F}}}\otimes{\mathcal{L}}(H),\quad{\mathcal{C}}({\mathbb{G}}):=\bigoplus_{\begin{subarray}{c}\alpha\in{\mathcal{F}}\\ 1\leq i,j\leq d_{\alpha}\end{subarray}}{\mathbb{C}}u^{\alpha}_{ij},$
the weak-norm Cauchy preservation assumed in the proof of Proposition˜1.6’s ˜dr $\Rightarrow$ ˜a implication holds for our $\psi_{\lambda}$ and the argument goes through. $\blacksquare$
@@
We will apply Lemma˜1.8, proving that its hypothesis obtains under the present circumstances.
Observe first that the individual positive operators $F_{\alpha_{n}}$ of ˜1-1 are identities: [32, Proposition 4.7] proves this globally, assuming the quantum group as a whole has property RD, but the argument (relying on [32, Lemmas 4.3 and 4.6]) goes through in the local version needed here.
Writing $H:=\bigoplus_{n}V_{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question, we have to argue that for $v\in H$ with
@@
($\operatorname{\mathrm{supp}}$ defined as in ˜1-8) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large. The unimodularity condition $F_{\alpha}=1$ ensures [35, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
@@
  • the free orthogonal group ${\mathbb{G}}:={\mathbb{O}}^{+}(n)$, $n\geq 3$ (underlying function algebra ${\mathcal{C}}({\mathbb{G}})=A_{o}(I_{n})$ in [32, §4.3]), with ${\mathcal{O}}({\mathbb{G}})$ freely generated by a unitary $n\times n$ matrix $u^{\alpha}=\left(u^{\alpha}_{ij}\right)$, $1\leq i,j\leq n$ of self-adjoint elements;
@@
will do by [32, Theorems 4.9 and 4.10] respectively. $\blacklozenge$